# supplement

### v2

@@ 4 Chebyshev filtered subspace iteration
Historically, subspace iteration has been overshadowed by Krylov methods, which are often more efficient when using standard arithmetic. However, in the context of low-rank approximation, subspace iteration offers distinct advantages over Krylov methods. As mentioned in the previous section, a major challenge in low-rank Krylov methods is that orthogonal Krylov basis vectors cannot be accurately represented in low-rank tensor formats, leading to significant truncation errors that hinder convergence. In contrast, subspace iteration methods can directly target low-rank eigenvectors as the basis for the subspace, which are orthonormal when $A$ is symmetric, bypassing the need for a Krylov basis. After discussing low-rank inexact subspace iteration algorithms, we demonstrate in the subsequent sections that the convergence of low-rank subspace iteration is more robust to truncation errors compared to Krylov methods. Additionally, we show that subspace iteration can efficiently compute several low-rank eigenvectors with machine precision.
@@ 4.1 Power iteration
To illustrate the rank truncated subspace iteration in the simplest setting, consider the power iteration – a subspace iteration involving a single vector. In power iteration, the matrix $A$ is repeatedly applied to an initial vector, and under certain conditions, this process converges to the dominant eigenvector of $A$. When $A$ and the vector being iterated are represented as low-rank tensors, applying $A$ to the current iterate typically results in a vector of higher rank. In rank-truncated power iteration, this vector is truncated and then normalized before proceeding to the next iteration. The main steps of power iteration with rank truncation are presented in Algorithm 2.
As we mentioned in Section 2.1.3 there are different approaches to computing the rank truncated matrix-vector product in line 3 of the algorithm. One approach is to represent the matrix-vector product using tensor cores with larger dimensions and then perform a TT-SVD truncation. This can become costly when the rank of $A$ or $v^{(i-1)}$ are moderately large, since under matrix-vector multiplication these ranks multiply. An alternative strategy is to approximate the matrix-vector product in the tangent space of the low-rank TT manifold at $v^{(i-1)}$ as described in Section 2.1.4. This approach does not require constructing intermediate tensors with large ranks and is therefore much more efficient when the rank of $A$ or $v$ is large.
@@
1:$v^{(0)}\in\mathbb{C}^{n_{1}\cdots n_{d}}$ low-rank TT
2:$v^{(i)}$ low-rank approximation to dominant eigenvector
3:$i=1$
4:while not converged do
5:     $z^{(i)}=\mathfrak{T}_{\bm{r}}\left(Av^{(i-1)}\right)$
6:     $v^{(i)}=z^{(i)}/\|z^{(i)}\|$
7:     $i=i+1$
8:end while
@@ 4.2 Subspace iteration
A basic form of the subspace iteration begins with a $m$ dimensional basis. One iteration involves applying the matrix $A$ to each basis vector and then orthonormalizing the resulting vectors against each other. The matrix multiplication amplifies components of each vector in the dominant eigendirection and the orthonormalization ensures that each vector converges to a different eigenvector. Under suitable conditions, iterating these steps yields a sequence of bases that eventually span the dominant eigenspace of $A$. We consider a low-rank variant of subspace iteration where all vectors are approximated using low-rank TTs. To describe one iteration of the low-rank TT subspace iteration suppose we have a basis of $m$ rank-$\bm{r}$ TTs $(v_{1},\ldots,v_{m})$. To perform one iteration we apply the matrix $A$ to each basis vector and truncate the result
(12) $z_{j}=\mathfrak{T}_{r}\left(Av_{j}\right),\qquad j=1,2,\ldots,m.$
Just as with the power iteration described in Section 4.1, different strategies can be used to approximate the matrix-vector products in (12) with low-rank TTs, such as TT-SVD or tangent space projection. Then, we construct new basis vectors by (approximately) orthonoromalizing the $z_{j}$ in (12) using a low-rank variant of the Rayleigh-Ritz projection. The Rayleigh-Ritz procedure is applied to ensure that the updated basis vectors are Ritz vectors, providing approximations of the dominant eigenvectors of $A$, which we assume can be represented as low-rank TTs. Note that in subspace iteration implementations without rank truncation, it is typical to use a different orthogonalization procedure such as QR-decomposition on the vectors (12). This however faces the same issue as Krylov subspace methods. In general, the orthogonal basis vectors are not low-rank making them difficult to represent efficiently in the TT format. We summarize the main steps of low-rank subspace iteration in Algorithm 3. In the algorithm we denote by $\texttt{RR}(A,z_{1},\ldots,z_{m})$ a call to the low-rank Rayleigh-Ritz procedure described hereafter.
@@
1:$A\in\mathbb{C}^{(n_{1}\cdots n_{d})\times(n_{1}\cdots n_{d})}$ TT-matrix, $v_{1}^{(0)},\ldots,v_{m}^{(0)}\in\mathbb{C}^{n_{1}\cdots n_{d}}$ initial system of $m$ low-rank tensors
2:$v_{1}^{(i)},\ldots,v_{m}^{(i)},{\lambda}_{1}^{(i)},\ldots,{\lambda}_{m}^{(i)}$ approximations of the $m$ dominant eigenpairs of $A$.
3:$i=1$
4:while not converged do
5:     $z_{j}^{(i)}=\mathfrak{T}_{\bm{r}}\left(Av_{j}^{(i-1)}\right),\qquad j=1,\ldots,m$
6:     $(v_{1}^{(i)},\ldots,v_{m}^{(i)}),(\lambda_{1}^{(i)},\ldots,\lambda_{m}^{(i)})=\texttt{RR}(A,z_{1}^{(i)},\ldots,z_{m}^{(i)})$
7:     $i=i+1$
8:end while
@@ 4.3 Rayleigh-Ritz procedure
(13) $\quad W_{ij}=\left\langle z_{i},z_{j}\right\rangle,\quad P_{ij}=\left\langle z_{i},Az_{j}\right\rangle\qquad i,j=1,2,\ldots,m,$
@@
(15) $\tilde{v}_{j}=\mathfrak{T}_{r}\left(\sum_{i=1}^{m}\Phi_{ij}z_{i}\right),\qquad i=1,2,\ldots,m.$
@@
Algorithm 4 Rayleigh-Ritz projection with rank truncation
1:$A\in\mathbb{C}^{(n_{1}\cdots n_{d})\times(n_{1}\cdots n_{d})}$, $(z_{1},\ldots,z_{m})$ low-rank basis vectors spanning a subspace of $\mathbb{C}^{n_{1}\cdots n_{d}}$.
2:$(v_{j},\lambda_{j})_{j=1}^{m}$ approximate eigenpairs of $A$.
3:$W_{jk}=\left\langle z_{j},z_{k}\right\rangle,\qquad j,k=1\ldots,m$
4:$P_{jk}=\left\langle z_{j},Az_{k}\right\rangle,\qquad j,k=1\ldots,m$
5:$\Phi,{\Lambda}$leftarrow $P\Phi={\Lambda}W\Phi$triangleright solve generalized eigenvalue problem
6:$\tilde{v}_{k}=\mathfrak{T}_{\bm{r}}\left(\sum_{j=1}^{m}\Phi_{jk}z_{j}\right),\qquad k=1,\ldots,m$
7:${v}_{j}=\tilde{v}_{j}/\left\|\tilde{v}_{j}\right\|,\qquad j=1,\ldots,m$
@@ 4.4 Chebyshev polynomial filtering
(16) $c_{k}(t)=\begin{cases}\cos\left(k\cos^{-1}(t)\right),&t\in[-1,1],\\ \cosh\left(k\cosh^{-1}(t)\right),&t\not\in[-1,1],\end{cases}$
@@
The polynomials $c_{k}$ are bounded inside of the interval $[-1,1]$ and rapidly increase outside $[-1,1]$. To construct a polynomial $p$ that yields a small ratio $p(\lambda_{m+1})/p(\lambda_{j})$, we compose $c_{k}$ with a linear transformation that maps $\lambda_{1},\ldots,\lambda_{m}$ outside of $[-1,1]$ and $\lambda_{m+1},\ldots,\lambda_{n_{1}\cdots n_{d}}$ inside $[-1,1]$. To construct such a linear map, suppose that all eigenvalues of $A$ lie within the real interval $[a_{0},b]$ and the eigenstates of interest $\lambda_{1},\ldots\lambda_{m}$ lie in the interval $[a_{0},a]$. The linear transformation
(18) $l(t)=\frac{t-c}{e},\qquad c=\frac{a+b}{2},\qquad e=\frac{b-a}{2},$
@@
maps $[a,b]$ into $[-1,1]$ and $[a_{0},a)$ outside of $[-1,1]$. Therefore, the polynomial $p_{k}=c_{k}\circ l$ is bounded inside $[a,b]$ and rapidly increases in $[a_{0},a)$, resulting in a large eigenvalue ratio $p_{k}(\lambda_{m+1})/p_{k}(\lambda_{j})$. In practice the interval $[a,b]$ is not known a priori but can be estimated. For example, we can set $b$ to the largest eigenvalue (or its approximation), and the endpoint $a$ can be updated after each Rayleigh-Ritz step to the largest Ritz value.
@@
To perform low-rank Chebyshev filtered power (or subspace) iteration we replace line 3 in Algorithm 2 (or Algorithm 3 for subspace) with a low-rank approximation of $p_{k}(A)v$. To compute this low-rank approximation, we utilize the three-term recurrence (17), inserting rank truncations to control the intermediate tensor ranks. This yields the rank-truncated three-term recurrence
(19) $q_{0}=v,\quad q_{1}=\mathfrak{T}_{\bm{r}}\left(\frac{Av-cv}{e}\right),\quad q_{j+1}=\mathfrak{T}_{\bm{r}}\left(\frac{2\mathfrak{T}_{\bm{r}}\left(Aq_{j}-cq_{j}\right)}{e}-q_{j-1}\right),$
where $q_{k}$ is a rank-$\bm{r}$ approximation of $p_{k}(A)v$. As mentioned above, there are different approaches to computing low-rank approximations of matrix-vector products and sums of TTs. Using TT-SVD requires applying multiple truncation procedures as written in (19). In contrast, tangent space projections or variational methods allow a single step of the three-term recurrence to be directly approximated with a low-rank TT, often leading to more efficient computations.
5 Convergence
In this section, we discuss of the convergence behavior of the Lanczos method and subspace iteration with rank truncation.
5.1 Lanczos
(20) $V_{m}=\hat{V}_{m}+E_{m},$
Projecting $A$ into an orthonormal basis of $V_{m}$ yields a $m\times m$ projected matrix that can be viewed as a perturbation of the tridiagonal matrix $T_{m}$ in (9) generated by truncation-free Lanczos iteraion. However, since the perturbation can be large, the eigenvalues of such a projected matrix (the Ritz values relative to the inexact Krylov subspace) can be significantly different from the eigenvalues of $T_{m}$ (the Ritz values of the Krylov subspace). Moreover, a Ritz vector obtained from the approximate Lanczos procedure cannot be expressed as $f_{m-1}(A)v_{0}$ for some polynomial $f_{m-1}(\lambda)$ of degree $m-1$. Consequently, the accuracy of the approximation cannot be readily analyzed using the polynomial approximation theory typically employed to study the convergence of the Lanczos algorithm [27]. It is, however, well-known that large errors during the early iterations of the Lanczos method, such as those introduced by rank truncation, can lead to stagnation in convergence [25]. This stagnation is observed in our numerical experiments presented in Section 7.
@@ 5.2 Power iteration
(21) $\mathcal{P}=\psi_{1}\psi_{1}^{\text{T}},\qquad\mathcal{P}_{\perp}=I_{n_{1}\cdots n_{d}}-\mathcal{P}.$
@@
(22) $\tan\measuredangle(Av,\psi_{1})\leq\left|\frac{\lambda_{2}}{\lambda_{1}}\right|\tan\measuredangle(v,\psi_{1}).$
For completion, we provide a proof of this theorem in Appendix A. The inequality (22) stated in the theorem can be visualized by projecting the vector space $\mathbb{C}^{n_{1}\cdots n_{d}}$ onto $\mathbb{R}^{2}$, where the first component represents the norm of the vector’s component in the range of $\mathcal{P}$ (i.e., in the direction of $\psi_{1}$), while the second component represents the norm of the vector’s component in the range of $\mathcal{P}_{\perp}$ (i.e., in the subspace orthogonal to $\psi_{1}$). We denote such projected vectors using a hat, e.g.,
(23) $\widehat{Av}=\begin{bmatrix}\|\mathcal{P}Av\|\\ \|\mathcal{P}_{\perp}Av\|\end{bmatrix}\quad\text{and}\quad\widehat{v}=\begin{bmatrix}\|\mathcal{P}v\|\\ \|\mathcal{P}_{\perp}v\|\end{bmatrix}.$
@@ 5.2.1 Single power iteration with truncation
(24) $z=Av+e,$
where $e$ is a perturbation due to rank truncation and we suppressed the iteration index $i$ while examining a single iteration. Progress towards convergence is achieved in a single iteration if the angle between $z$ and $\psi_{1}$ is smaller than the angle between $v$ and $\psi_{1}$, i.e., $\measuredangle(z,\psi_{1})<\measuredangle(v,\psi_{1})$. Such progress is guaranteed when the perturbation $e$ introduced by rank-truncation is sufficiently small. The acceptable size of $e$ that guarantees convergence is illustrated Fig. 1(a). As long as the truncated vector remains within the green circle, the condition $\measuredangle(z,\psi_{1})<\measuredangle(v,\psi_{1})$ holds, ensuring progress towards convergence. Hence the radius of the green circle is the maximum allowable truncation error that ensures convergence for perturbations $e$ in arbitrary directions. This radius is the norm of the component of $\widehat{Av}$ that is orthogonal to $v$, which provides an upper bound for the truncation error
(25) $\|e\|<\left\|\left(I_{2}-\widehat{v}\widehat{v}^{\text{T}}\right)\widehat{Av}\right\|,$
@@
(26) $\frac{\|e\|}{|\lambda_{1}|}<\left(1-\left|\frac{\lambda_{2}}{\lambda_{1}}\right|\right)\frac{\cos(\theta)\sin(\theta)}{\cos(\theta)+\sin(\theta)},$
@@
Proof 5.3.
Substituting $\cos(\theta)=\|\mathcal{P}v\|$ and $\sin(\theta)=\|\mathcal{P}_{\perp}v\|$ we rewrite the inequality (26) as
(27) $\|e\|<\left(|\lambda_{1}|-|\lambda_{2}|\right)\frac{\|\mathcal{P}v\|\cdot\|\mathcal{P}_{\perp}v\|}{\|\mathcal{P}v\|+\|\mathcal{P}_{\perp}v\|}.$
@@
(28) $\frac{|\lambda_{1}|\|\mathcal{P}v\|-\|e\|}{|\lambda_{2}|\|\mathcal{P}_{\perp}v\|+\|e\|}>\frac{\|\mathcal{P}v\|}{\|\mathcal{P}_{\perp}v\|}.$
It follows from (56) that the left hand side of (28) can be bounded from above by
(29) $\frac{\|\mathcal{P}Av\|-\|e\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\leq\frac{\|\mathcal{P}Av\|-\|\mathcal{P}e\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\leq\frac{\|\mathcal{P}Av+\mathcal{P}e\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\\ \leq\frac{\|\mathcal{P}\left(Av+e\right)\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}$
@@
(30) $\frac{\|\mathcal{P}\left(Av+e\right)\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\geq\frac{|\lambda_{1}|\|\mathcal{P}v\|-\|e\|}{|\lambda_{2}|\|\mathcal{P}_{\perp}v\|+\|e\|}>\frac{\|\mathcal{P}v\|}{\|\mathcal{P}_{\perp}v\|},$
completing the proof.
@@
(31) $\theta^{(i)}=\measuredangle\left(v^{(i)},\psi_{1}\right).$
@@
(32) $\sin\left(\theta^{(i)}\right)=\left\|\mathcal{P}_{\perp}v^{(i)}\right\|,$
@@
(33) $\left\|\mathfrak{T}_{\bm{r}}\left(v^{(i)}\right)-v^{(i)}\right\|\leq\left\|\mathcal{P}_{\perp}v^{(i)}\right\|=\sin\left(\theta^{(i)}\right).$

### v3

@@ 3 Chebyshev filtered subspace iteration
Historically, subspace iteration has been overshadowed by Krylov methods, which are often more efficient when using standard arithmetic. However, in the context of low-rank approximation, subspace iteration offers distinct advantages over Krylov methods such as those proposed in [5, 19, 20, 26] and summarized in section A.1. A major challenge in low-rank Krylov methods is that orthogonal Krylov basis vectors cannot be accurately represented in low-rank tensor formats, leading to significant truncation errors that hinder convergence. In contrast, subspace iteration methods can directly target low-rank eigenvectors as the basis for the subspace bypassing the need for a Krylov basis. After discussing low-rank inexact subspace iteration algorithms, we demonstrate in the subsequent sections that the convergence of low-rank subspace iteration is more robust to truncation errors compared to Krylov methods.
@@ 3.1 Power iteration
To illustrate the rank truncated subspace iteration in the simplest setting, consider the power iteration – a subspace iteration involving a single vector. In power iteration, the matrix $A$ is repeatedly applied to an initial vector, and under certain conditions, this process converges to the dominant eigenvector of $A$. When $A$ and the vector being iterated are represented as low-rank tensors, applying $A$ to the current iterate typically results in a vector of higher rank. In rank-truncated power iteration, this vector is truncated to a user-provided rank $\bm{r}$ and then normalized before proceeding to the next iteration. The main steps of power iteration with rank truncation are presented in algorithm 1. In this algorithm, the operator $\mathfrak{T}_{\bm{r}}$ denotes an operator that truncates a TT to rank $\bm{r}$. The truncation rank $\bm{r}$ and the algorithm used to perform the truncation is set by the user. More details on choices of rank-truncation algorithms and their computational costs are discussed in section 5.1.
@@
1:
[…]
5:
[…]
8:repeat
9:  $z^{(i+1)}\leftarrow\mathfrak{T}_{\bm{r}}\!\left(Av^{(i)}\right)$ $\triangleright$ Apply AA and truncate
10:  $v^{(i+1)}\leftarrow z^{(i+1)}/\lVert z^{(i+1)}\rVert$ $\triangleright$ Normalize
11:  $i\leftarrow i+1$
12:until convergence
@@ 3.2 Subspace iteration
A basic form of the subspace iteration begins with a $m$ dimensional basis. One iteration involves applying the matrix $A$ to each basis vector and then orthonormalizing the resulting vectors against each other. The matrix multiplication amplifies components of each vector in the dominant eigendirection and the orthonormalization ensures that each vector converges to a different eigenvector. Under suitable conditions, iterating these steps yields a sequence of bases that eventually span the dominant eigenspace of $A$. We consider a low-rank variant of subspace iteration where all vectors are approximated using low-rank TTs. To describe one iteration of the low-rank TT subspace iteration suppose we have a basis of $m$ rank-$\bm{r}$ $T$s $(v_{1},\ldots,v_{m})$. To perform one iteration we apply the matrix $A$ to each basis vector and truncate the result
$z_{j}=\mathfrak{T}_{\bm{r}}\left(Av_{j}\right),\qquad j=1,2,\ldots,m.$ (6)
Just as with the power iteration described in section 3.1, the truncation rank $\bm{r}$ and the algorithm used to perform the truncation is set by the user. More details on choices of rank-truncation algorithms and their computational costs are provided in section 5.1. Then, we construct new basis vectors by approximately orthonoromalizing the $z_{j}$ in (6) using a low-rank variant of the Rayleigh-Ritz projection. The Rayleigh-Ritz procedure is applied to ensure that the updated basis vectors are Ritz vectors, providing approximations of the dominant eigenvectors of $A$, which we assume can be represented as low-rank TTs. Note that in subspace iteration implementations without rank truncation, it is typical to use a different orthogonalization procedure such as QR-decomposition on the vectors (6). This however faces the same issue as Krylov subspace methods summarized in appendix A. In general, the orthogonal basis vectors are not low-rank making them difficult to represent efficiently in the TT format. We summarize the main steps of low-rank subspace iteration in algorithm 2. In the algorithm we denote by $\texttt{RR}(A,z_{1},\ldots,z_{m})$ a call to the low-rank Rayleigh-Ritz procedure described hereafter.
@@
1:
[…]
5:
[…]
9:while not converged do
10:  $z_{j}^{(i)}\leftarrow\mathfrak{T}_{\bm{r}}\!\left(Av_{j}^{(i-1)}\right)$  for $j=1,\ldots,m$
11:  $(v_{1}^{(i)},\ldots,v_{m}^{(i)}),\,(\lambda_{1}^{(i)},\ldots,\lambda_{m}^{(i)})\leftarrow\texttt{RR}(A,z_{1}^{(i)},\ldots,z_{m}^{(i)})$
12:  $i\leftarrow i+1$
13:end while
@@ 3.3 Rayleigh-Ritz procedure
$\quad W_{ij}=\left\langle z_{i},z_{j}\right\rangle,\quad P_{ij}=\left\langle z_{i},Az_{j}\right\rangle\qquad i,j=1,2,\ldots,m,$ (7)
@@
$\tilde{v}_{j}=\mathfrak{T}_{r}\left(\sum_{i=1}^{m}\Phi_{ij}z_{i}\right),\qquad i=1,2,\ldots,m.$ (9)
@@
Algorithm 3 Rayleigh–Ritz projection with rank truncation
1:
[…]
5:
[…]
10:Solve $P\Phi=\Lambda W\Phi$ $\triangleright$ Generalized eigenproblem for Ritz values/vectors
[…]
@@ 3.4 Chebyshev polynomial filtering
$c_{k}(t)=\begin{cases}\cos\left(k\cos^{-1}(t)\right),&t\in[-1,1],\\ \cosh\left(k\cosh^{-1}(t)\right),&t\not\in[-1,1],\end{cases}$ (10)
@@
The polynomials $c_{k}$ are bounded inside of the interval [1,1][-1,1] and rapidly increase outside [1,1][-1,1]. To construct a polynomial $p$ that yields a small ratio $p(\lambda_{m+1})/p(\lambda_{j})$, we compose $c_{k}$ with a linear transformation that maps $\lambda_{1},\ldots,\lambda_{m}$ outside of [1,1][-1,1] and $\lambda_{m+1},\ldots,\lambda_{n_{1}\cdots n_{d}}$ inside [1,1][-1,1]. To construct such a linear map, suppose that all eigenvalues of $A$ lie within the real interval $a_{0},b]$ and the eigenstates of interest $\lambda_{1},\ldots\lambda_{m}$ lie in the interval $a_{0},a]$. The linear transformation
$l(t)=\frac{t-c}{e},\qquad c=\frac{a+b}{2},\qquad e=\frac{b-a}{2},$ (12)
@@
maps $[a,b]$ into [1,1][-1,1] and $a_{0},a)$ outside of [1,1][-1,1]. Therefore, the polynomial $p_{k}=c_{k}\circ l$ is bounded inside $[a,b]$ and rapidly increases in $a_{0},a)$, resulting in a large eigenvalue ratio $p_{k}(\lambda_{m+1})/p_{k}(\lambda_{j})$. In practice the interval $[a,b]$ is not known a priori but can be estimated. For example, we can set $b$ to the largest eigenvalue (or its approximation), and the endpoint $a$ can be updated after each Rayleigh-Ritz step to the largest Ritz value.
@@
To perform low-rank Chebyshev filtered power (or subspace) iteration we $Av$ with a low-rank approximation of $p_{k}(A)v$, where the polynomial filter degree $k$ is a user-supplied parameter. To compute this low-rank approximation, we utilize the three-term recurrence (11), inserting rank truncations to control the intermediate tensor ranks. This yields the rank-truncated three-term recurrence
$q_{0}=v,\quad q_{1}=\mathfrak{T}_{\bm{r}}\left(\frac{Av-cv}{e}\right),\quad q_{j+1}=\mathfrak{T}_{\bm{r}}\left(\frac{2\mathfrak{T}_{\bm{r}}\left(Aq_{j}-cq_{j}\right)}{e}-q_{j-1}\right),$ (13)
where $q_{k}$ is a rank-$\bm{r}$ approximation of $p_{k}(A)v$. As mentioned above, the operator $\mathfrak{T}_{\bm{r}}$ denotes an operator that truncates a TT to rank-$\bm{r}$. The truncation rank $\bm{r}$ and the algorithm used to perform the truncation is set by the user. More details on choices of rank-truncation algorithms and their computational costs are discussed in section 5.1. We summarize the Chebyshev filtered subspace iteration with rank truncation in algorithm 4.
Algorithm 4 Chebyshev filtered subspace iteration with rank truncation
1:
[…]
6:
[…]
10:while not converged do
11:  $z_{j}^{(i)}\leftarrow\mathfrak{T}_{\bm{r}}\!\left(p_{k}(A)v_{j}^{(i-1)}\right)$  for $j=1,\ldots,m$
12:  $(v_{1}^{(i)},\ldots,v_{m}^{(i)}),\,(\lambda_{1}^{(i)},\ldots,\lambda_{m}^{(i)})\leftarrow\texttt{RR}(A,z_{1}^{(i)},\ldots,z_{m}^{(i)})$
13:  $i\leftarrow i+1$
14:end while
4 Upper bounds on truncation error
In this section, we analyze the convergence behavior of TT subspace iteration with rank truncation. Unlike the analysis in [29], we derive an explicit upper bound on the inexactness at each iteration, which guarantees progress toward convergence. These results serve as bounds on the truncation error in rank-truncated power and subspace iterations that ensure convergence.
@@ 4.1 Power iteration
$\mathcal{P}=\psi_{1}\psi_{1}^{\text{T}},\qquad\mathcal{P}_{\perp}=I_{n_{1}\cdots n_{d}}-\mathcal{P}.$ (14)
@@
$\tan\measuredangle(Av,\psi_{1})\leq\left|\frac{\lambda_{2}}{\lambda_{1}}\right|\tan\measuredangle(v,\psi_{1}).$ (15)
The above theorem is a well-known result [27], and we provide a proof in appendix C. The inequality (15) stated in the theorem can be visualized by projecting the vector space $\mathbb{C}^{n_{1}\cdots n_{d}}$ onto $\mathbb{R}^{2}$, where the first component represents the norm of the vector’s component in the range of $\mathcal{P}$ (i.e., in the direction of $\psi_{1}$), while the second component represents the norm of the vector’s component in the range of $\mathcal{P}_{\perp}$ (i.e., in the subspace orthogonal to $\psi_{1}$). We denote such projected vectors using a hat, e.g.,
$\widehat{Av}=\begin{bmatrix}\|\mathcal{P}Av\|\\ \|\mathcal{P}_{\perp}Av\|\end{bmatrix}\quad\text{and}\quad\widehat{v}=\begin{bmatrix}\|\mathcal{P}v\|\\ \|\mathcal{P}_{\perp}v\|\end{bmatrix}$. (16)
@@ 4.1.1 Single power iteration with truncation
$z=Av+e,$ (17)
where $e$ is a perturbation due to rank truncation and we suppressed the iteration index $i$ while examining a single iteration. Progress towards convergence is achieved in a single iteration if the angle between $z$ and $\psi_{1}$ is smaller than the angle between $v$ and $\psi_{1}$, i.e., $\measuredangle(z,\psi_{1})<\measuredangle(v,\psi_{1})$. Such progress is guaranteed when the perturbation $e$ introduced by rank-truncation is sufficiently small. To determine the acceptable size of $e$ that guarantees convergence we consider the vectors (16) in fig. 1. As long as the truncated vector remains within the green circle, the condition $\measuredangle(z,\psi_{1})<\measuredangle(v,\psi_{1})$ holds, ensuring progress towards convergence. Hence the radius of the green circle is the maximum allowable truncation error that ensures convergence for perturbations $e$ in arbitrary directions. This radius is the norm of the component of $\widehat{Av}$ that is orthogonal to $v$, which provides an upper bound for the truncation error
$\|e\|<\left\|\left(I_{2}-\widehat{v}\widehat{v}^{\text{T}}\right)\widehat{Av}\right\|$, (18)
@@
$\frac{\|e\|}{|\lambda_{1}|}<\left(1-\left|\frac{\lambda_{2}}{\lambda_{1}}\right|\right)\frac{\cos(\theta)\sin(\theta)}{\cos(\theta)+\sin(\theta)},$ (19)
@@
Proof: Substituting $\cos(\theta)=\|\mathcal{P}v\|$ and $\sin(\theta)=\|\mathcal{P}_{\perp}v\|$ we rewrite the inequality (19) as
$\|e\|<\left(|\lambda_{1}|-|\lambda_{2}|\right)\frac{\|\mathcal{P}v\|\cdot\|\mathcal{P}_{\perp}v\|}{\|\mathcal{P}v\|+\|\mathcal{P}_{\perp}v\|}.$ (20)
@@
$\frac{|\lambda_{1}|\|\mathcal{P}v\|-\|e\|}{|\lambda_{2}|\|\mathcal{P}_{\perp}v\|+\|e\|}>\frac{\|\mathcal{P}v\|}{\|\mathcal{P}_{\perp}v\|}.$ (21)
It follows from
$\left\|\mathcal{P}_{\perp}Av\right\|\leq|\lambda_{2}|\left\|\mathcal{P}_{\perp}v\right\|\quad\text{and}\quad\left\|\mathcal{P}Av\right\|=|\lambda_{1}|\left\|\mathcal{P}v\right\|.$ (22)
that the left hand side of (21) can be bounded from above by
$\frac{\|\mathcal{P}Av\|-\|e\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\leq\frac{\|\mathcal{P}Av\|-\|\mathcal{P}e\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\leq\frac{\|\mathcal{P}Av+\mathcal{P}e\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\\ \leq\frac{\|\mathcal{P}\left(Av+e\right)\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}$ (23)
@@
$\frac{\|\mathcal{P}\left(Av+e\right)\|}{\left\|\mathcal{P}_{\perp}\left(Av+e\right)\right\|}\geq\frac{|\lambda_{1}|\|\mathcal{P}v\|-\|e\|}{|\lambda_{2}|\|\mathcal{P}_{\perp}v\|+\|e\|}>\frac{\|\mathcal{P}v\|}{\|\mathcal{P}_{\perp}v\|},$ (24)
completing the proof. $\square$
@@
$\theta^{(i)}=\measuredangle\left(v^{(i)},\psi_{1}\right).$ (25)
@@
$\sin\left(\theta^{(i)}\right)=\left\|\mathcal{P}_{\perp}v^{(i)}\right\|,$ (26)
@@
$\left\|\mathfrak{T}_{\bm{r}}\left(v^{(i)}\right)-v^{(i)}\right\|\leq\left\|\mathcal{P}_{\perp}v^{(i)}\right\|=\sin\left(\theta^{(i)}\right).$ (27)